A mechanic needs a radiator to have 40% antifreeze solutions. the radiator currently is filled with 4 gallons of 25%antifreeze solution. How much antifreeze mixture should be drained from the car if the mechanic replaces it with pure antifreeze.
Drain_____ gallons from the radiator(if needed answer to 3 decimal places)

Answer by addingup(3677)

If he drains x gallons, then
.25(4-x) + 1.00x = .40(4)
1 - .25x + x = 1.6
.75x = .6
x = .8
So, after draining .8 gals, he has 3.2 gals of 25% antifreeze. This means he has That means he has 3.2(0.25)= .8 gals of antifreeze in the radiator.
Now he adds .8 gals of pure antifreeze. Then he has 1.6 gals of antifreeze in 4 gals of mixture. That's .16/4 = 40% antifreeze. We've got the right answer.